OpenStudy (whiteduckling):

what's the area of a circle with a radius of 12?

OpenStudy (jasonnewstedrocks):

The formula is: A =πr^2

OpenStudy (astrophysics):

The area of a circle can be found using \[\huge A = \pi r^2\] where r is the radius, plug and chug!

OpenStudy (jasonnewstedrocks):

So first is 3.14*12

OpenStudy (anonymous):

It would be A= pie* R^2 So it would be A= 3.14* 12^2 Pie = 3.14

OpenStudy (whiteduckling):

just tell me

OpenStudy (jasonnewstedrocks):

Nooo.

OpenStudy (jasonnewstedrocks):

Com' on 3.14*12^2

OpenStudy (jasonnewstedrocks):

You can do 3.14*12 first, and then ^2

OpenStudy (whiteduckling):

452.16

OpenStudy (ineedhelplz):

A≈452.39 Your're right :)

OpenStudy (jasonnewstedrocks):

Ye

OpenStudy (whiteduckling):

ok thx